OpenStudy (anonymous):

a set of 24 cards in numbered with the positive integers from 1 to 24. if the cards are shuffled and if only one is selected at random, what is the probability that the number on the card is divisible by 4 or 6?

OpenStudy (mathmate):

How many of the numbers between 1 and 24 are divisible by 4 or by 6? Divide this by the size of the sample space (24) to get the probability.

OpenStudy (anonymous):

divided by 4 are 4,8,12,16,20,24 6 are 6,12,18,24 10/24=5/12 is it correct??

OpenStudy (anonymous):

@mathmate

OpenStudy (raden):

the question is little ambigu,which one the correct : divisible by 4 or divisible by 6 ? or divisible by 4 or 6 ? --> we knowed that the number of 6 only once here

ganeshie8 (ganeshie8):

haha true its ambiguous ^

ganeshie8 (ganeshie8):

divided by 4 are 4,8,12,16,20,24 6 are 6,12,18,24 10/24=5/12 is it correct??

ganeshie8 (ganeshie8):

@dinisha - careful, there are some common numbers between the two sets

ganeshie8 (ganeshie8):

divided by 4 are 4,8, `12`,16,20, `24` 6 are 6, `12`,18, `24`

OpenStudy (mathmate):

I think the 'OR' is just the normal interpretation, not the exclusive or. It would otherwise have been qualified as "but not both".

OpenStudy (anonymous):

oh so when there is repeatition it is only count once? @ganeshie8

ganeshie8 (ganeshie8):

Exactly !

ganeshie8 (ganeshie8):

P(divisible by 4) = 6/24 P(divisible by 6) = 4/24 P(divisible by 4 and 6) = 2/24 So, P(divisible by 4 or 6) = 6/24 + 4/24 - 2/24 = 8/24 = 1/3

OpenStudy (anonymous):

thank you.. i am clear now!!

ganeshie8 (ganeshie8):

good :)